OpenStudy (anonymous):

What is 1 - secant? I couldnt find that anywhere, its in my textbook. I triple checked, its 1 - secant not secant - 1. Really need help, what is 1 - secant of theta?

OpenStudy (anonymous):

\[\sec \theta=1/\cos \theta \]

OpenStudy (anonymous):